OpenStudy (anonymous):

The volumes of two similar solids are 8ft^2 and 125ft^2. The surface area of the smaller solid is 4ft^2 . What is the surface area of the larger solid? A) 25 ft^2 B) 32ft^2 C) 50ft^2 D) 16ft^2

OpenStudy (anonymous):

\[\sqrt[3]{V_1/V_2} = \sqrt{A_1/A_2}\] \[\sqrt[3]{8/125} = \sqrt{4/x}\] \[2/5 = 2/ \sqrt{x}\] \[x = 25\] Answer: A)